1) g(x)=5^x Find the domain .
2) g(x)=e^x-5 Describe the transformations:
equation(s) for the vertical asymptote(s):
x-intercept in (x,y) form:
3)g(x)=log(x-3) same for #2
4)P=11R44LN x where x is the # of years after 2002

Answer by stanbon(75887)

1) g(x)=5^x Find the domain------------All Real Numbers
---------------------------------------------------- .
2) g(x)=e^(x-5) Describe the transformations: shift h(x)=e^x five steps to the right.
equation(s) for the vertical asymptote(s): no vertical asymptote
x-intercept in (x,y) form: if y = 0, e^(x-5) = 0; no solution because y = 0
is the horizontal asymptote of g(x)
---------------------------------------------------------
3)g(x)=log(x-3) same for #2
shift h(x) = logx three to the right
--------
vertical asymptote at x = 3
--------
4)P=11R44LN x where x is the # of years after 2002
What is the question?
=============================
Cheers,
Stan H.